OpenStudy (anonymous):

Use the quadratic formula to find the solution to the quadratic equation given below. x^2-5x+25/4=0

OpenStudy (mathstudent55):

Is this the equation? \( x^2 - 5x + \dfrac{25}{4} = 0\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x ^{2}-5 x+25/4=0,or 4 x ^{2}-20 x+25=0\] a=4,b=-20,c=25 \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\] solve

OpenStudy (mathstudent55):

Here you have a = 1, b = -5, and c = 25/4. The quadratic formula is: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) We substitute our values of a, b, and c, and then we simplify. \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) \(x = \dfrac{-(-5) \pm \sqrt{(-5)^2 - 4(1)\frac{25}{4}}}{2(1)} \) \(x = \dfrac{5 \pm \sqrt{25 - 25}}{2} \) \(x = \dfrac{5 \pm \sqrt{0}}{2} \) \(x = \dfrac{5}{2} \)

OpenStudy (anonymous):

it should be ( -20 )^{2}=400 & not -20^{2}

OpenStudy (mathstudent55):

If you prefer, you can do it this way: \(4x^2 - 20x + 25 = 0\) with a = 4, b = -20, c = 25 \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) \(x = \dfrac{20 \pm \sqrt{400 - 4(4)(25)}}{2(4)}\) \(x = \dfrac{20}{8} = \dfrac{5}{2}\) The result is the same.